# Make a Donut: Hierarchical EMD-Space Planning for Zero-Shot Deformable Manipulation with Tools

### v2

@@ V Conclusions and Limitations
We introduced a new hierarchical planning method for robotic deformable object manipulation, enabling complex tasks without prior demonstrations. This method surpasses previous demonstration-based techniques, ensuring better adaptability to new scenarios. Using large language models, it generates high-level plans and intermediate goals, which are executed through a unique closed-loop predictive control using Differentiable Physics. Our approach showed exceptional performance and adaptability in dough manipulation benchmarks, marking a significant step forward in deformable object manipulation.
@@
While our method is adept at handling complex tasks involving long-horizon planning, it is limited to generating simple shapes that can be described with the Python codes produced by the LLM. Generating Python code to describe the shape of an arbitrary object is extremely challenging, if not impossible, for nowaday’s LLM. However, we have made efforts to circumvent this limitation by having the LLM generate intermediate text descriptions that can be input into state-of-the-art Text-to-3D generative models, such as Point-E [31]. Figure 9 illustrates an example of creating a bowl from the dough. The shape of a bowl is difficult to describe using pure Python code, so the LLM outputs text descriptions for each subgoal, like “Flattened Sphere” and “Bowl”. These textx are then interpreted by Point-E to generate point clouds. We then proceed with our planning, using these point clouds as intermediate targets.
Besides, the dimensions and morphology of candidate tools are pre-determined as a priori constraints, consequently limiting the range of tasks that can be undertaken. A compelling avenue for future research could be to investigate the simultaneous optimization of the size of these tools, or let the LLM be aware/decide the size of the tools.
@@ A-B Hyperparameters for Simulation Dough
We use PlasticineLab [5] to simulate the dough. PlasticineLab is a Python environment that is suitable for deformable object simulation. The hyperparameters that are relevant to the properties of the dough are identical to those in PASTA [4] to ensure a fair comparison.
@@ A-D Intermediate Subgoal Visualization
Figure 11 depicts more execution trajectories of our method, including intermediate goals generated by the Language Learning Model (LLM). While LLM may not produce subgoals that are perfectly reachable for execution, the closed-loop execution with EMD space point-to-point planning (DiffPhysics-P2P) algorithm within our approach exhibits resilience to such imperfections.
@@ A-E Next Reachable Point Cloud Visualization
In this section, we visualize the topological change of next reachable point clouds throughout the process of making a donut. The visualizations demonstrate key stages, such as the intentional alteration of the dough’s topology to create a hole in the middle of a donut. In this context, the intrinsic property of optimal transport (OT) that does not necessarily preserve topological structures is leveraged to our advantage. Although our tools may not execute the point cloud to perfection, the LLM-generated subgoals guide the gradient flow towards the target. The methodology allows for self-correction with each observation, ensuring that each step brings us closer to the intended outcome.
Figure 12: Next reachable point cloud visualization for making a donut. Notice that the topology is gradually changed during the second stage for making the donut. The LLM planning phase ensures that subsequent OT steps are guided towards physically realizable subgoals.

### v3

@@ V Conclusions and Limitations
We introduced a new hierarchical planning method for robotic deformable object manipulation, enabling complex tasks without prior demonstrations. Using large language models, it generates high-level plans and intermediate goals, which are executed through a unique closed-loop predictive control using Differentiable Physics. Our approach showed better performance and adaptability in dough manipulation benchmarks, marking a significant step forward in deformable object manipulation.
@@
While our method is adept at handling complex tasks involving long-horizon planning, it is limited to generating simple shapes that can be described with the Python codes produced by the LLM. However, intermediate text descriptions can be converted by Text-to-3D generative models, such as Point-E [31]. Figure 9 illustrates an example of creating a bowl from the dough. The shape of a bowl is difficult to describe using pure Python code, so the LLM outputs text descriptions for each subgoal, like Flattened Sphere and Bowl. These texts are then interpreted by Point-E to generate point clouds. We then proceed with our planning, using these point clouds.
@@ A-B Hyperparameters for Simulation Dough
We use PlasticineLab [5] to simulate the dough. The physical parameters of the dough are set identical to those in PASTA [4] to ensure a fair comparison.
@@ A-D Intermediate Subgoal Visualization
Figure 11 depicts more execution trajectories of our method, including intermediate goals generated by the LLM. While LLM may not produce subgoals that are perfectly reachable for execution, the closed-loop execution with EMD space point-to-point planning (DiffPhysics-P2P) algorithm within our approach exhibits resilience to such imperfections.
@@ A-E Next Reachable Point Cloud Visualization
In this section, we visualize the topological change of next reachable point clouds throughout the process of making a donut. The visualizations demonstrate key stages, such as the intentional alteration of the dough’s topology to create a hole in the middle of a donut. Although our tools may not execute the point cloud to perfection, the LLM-generated subgoals guide the gradient flow towards the correct target.
Figure 12: Next reachable point cloud visualization. Notice that the topology is gradually changed during the second stage for making the donut. LLM planning ensures that subsequent steps are guided towards physically realizable subgoals.